OpenStudy (anonymous):

I NEED HELP ASAP. I AM CURRENTLY 6 LESSON DAYS BEHIND AND VERY DEPRESSED/STRESSED!! You are making a rectangular table. The area of the table should be 10 ft^2. You want the length of the table to be 1 ft shorter than twice its width. What should the dimensions of the table be? Okay, so, I'm not sure how to put this into an equation. The example in the book is very different to this problem.

jimthompson5910 (jim_thompson5910):

let w = width "You want the length of the table to be 1 ft shorter than twice its width" so the length is 2w-1

jimthompson5910 (jim_thompson5910):

Area = Length * Width A = L*W 10 = (2w-1)*w 10 = w*(2w-1) 10 = 2w^2 - w agreed so far?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

get everything to one side 10 = 2w^2 - w 10-10 = 2w^2 - w - 10 0 = 2w^2 - w - 10 2w^2 - w - 10 = 0

jimthompson5910 (jim_thompson5910):

to solve 2w^2 - w - 10 = 0, you can factor but using the quadratic formula is the best option

OpenStudy (anonymous):

Yeah, we're currently using quadratic formulas.

jimthompson5910 (jim_thompson5910):

2w^2 - w - 10 = 0 means a = 2 b = -1 c = -10 plug those values of a,b,c into the quadratic formula

OpenStudy (anonymous):

2x^2+-x+-10=0 Like that?

jimthompson5910 (jim_thompson5910):

yes you can swap each w for x

jimthompson5910 (jim_thompson5910):

and rewrite it like that, yep

OpenStudy (anonymous):

Okay, thank you very much! I've pretty much got everything from here. ;D

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

Let me show you how to find the solutions using the quadratic formula

jimthompson5910 (jim_thompson5910):

Quadratic Formula: \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

a = 2 b = -1 c = -10 \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-(-1) \pm \sqrt{(-1)^2-4*(2)*(-10)}}{2*(2)}\] \[\Large x=\frac{1 \pm \sqrt{81}}{4}\] \[\Large x=\frac{1 + \sqrt{81}}{4} \text{ or } x=\frac{1 - \sqrt{81}}{4}\] \[\Large x=\frac{1 + 9}{4} \text{ or } x=\frac{1 - 9}{-2}\] \[\Large x=\frac{10}{4} \text{ or } x=\frac{-8}{4}\] \[\Large x=\frac{5}{2} \text{ or } x=-2\] \[\Large x=2.5 \text{ or } x=-2\]

OpenStudy (anonymous):

Yeah, that is what the algebra calculators said too but for some reason, the back of the book says the answer is 2.5 ft by 4 ft, which is what makes me so confused.

jimthompson5910 (jim_thompson5910):

well x = -2 is tossed out because we can't have a negative width

jimthompson5910 (jim_thompson5910):

x = 2.5 is the only possible width

jimthompson5910 (jim_thompson5910):

x = width ""You want the length of the table to be 1 ft shorter than twice its width"" so length = 2x - 1

jimthompson5910 (jim_thompson5910):

length = 2x - 1 length = 2*2.5 - 1 length = 4

OpenStudy (anonymous):

Ooh, thanks so much!! ;'D